Discuss the Seebeck effect and one of its applications.


1
Expert's answer
2021-07-21T09:06:13-0400

Answer: Seebeck effect - the phenomenon of electromotive force between two contacts of different conductors, which are at different temperatures. An example of an application is a thermocouple.

The phenomenon of stress in a metal bar, the ends of which had different temperatures, was discovered (accidentally) in 1821 by Thomas Seebeck.

The Seebeck effect occurs in a circuit consisting of two soldered conductors (thermocouple). One of the contacts is heated, and then an electric current occurs in the circuit. The potential difference (electromotive force) that occurs between the contacts depends on the type of contact conductors and the temperature difference between the contacts. It can be presented in the form of a general formula:

"E=f(T_1)-f(T_2), T_1>T_2" .

The derivative of the function f (T)

"\alpha=\dfrac{df(T)}{dT}"

is called the coefficient of thermoelectromotive force of contact.

The coefficient of thermoelectromotive force of the contact can be given as the difference of two values ​​that characterize the material of the conductor:

"\alpha_{12}=\alpha_1-\alpha_2"

The coefficients "\alpha_i" are called Seebeck coefficients or coefficients of thermo-driving force of matter. In general, these coefficients depend on the temperature, and may even change sign as it changes.